# Detecting AI-Generated Text: Factors Influencing Detectability with Current Methods

### v1

@@ 1 Introduction
In recent years, the capabilities of large language models (LLMs) to generate fluent, realistic-sounding text have improved dramatically. We are now at a point where humans themselves cannot reliably distinguish between text that was generated using artificial intelligence (AI) and that written by a real person (?, ?, ?). There are many opportunities for LLMs to contribute to human productivity, with applications to question-answering, computer programming, brainstorming, proof-reading, and information retrieval. However, there are also potential areas for misuse, such as academic dishonesty, professional misconduct, and malicious chatbots (?, ?). Another major risk brought by generative AI is the potential for global-scale “information pollution.” This includes disinformation, where false information is knowingly shared with the intent of causing harm, as well as misinformation, where the false information is shared unknowingly (?). These categories encompass a wide range of attacks including propaganda, rumours, ‘fake news’, fraud, etc. Current state-of-the-art generative models can produce high-quality, fluent fake information that is perceived as more credible and trustworthy than human-generated misinformation (?, ?), and that is harder for both human readers and automatic detection systems to recognize (?, ?, ?). Figure 1 outlines a few of the risks of widely available AI text generation tools.
Figure 1: High-level overview of the risks posed by AI-generated text (?, ?).
@@
The present manuscript provides an overview of the field of AIGT detection, the current state-of-the-art methodologies, the available data resources and online tools, and existing challenges. In this survey, we have included papers published in both peer-reviewed journals and conference proceedings, as well as those posted on pre-print servers such as arXiv. While pre-print manuscripts may not have been formally peer-reviewed yet, we include them in an attempt to cover the most cutting-edge advancements in a rapidly changing field. We have focused our attention primarily on papers published in 2023 and 2024, with some references to seminal papers published earlier. We include only those papers that focus on NLP methods for AIGT detection, excluding methods that rely on other characteristics of bot accounts, such as posting frequency or social network analysis. However, within the realm of NLP, we cast our net widely, including methods from diverse data domains (social media, news stories, academic essays, scientific abstracts) and aiming to provide an extensive and high-level overview of the field as it currently stands.
@@
Other surveys on AIGT detection exist, and we point the interested reader to them for more information. ? (?) and ? (?) focus specifically on watermarking for LLMs. ? (?) provide an extensive analysis of the risks of AI-generated text, and an evaluation of methods to detect text generated by pre-ChatGPT models. ? (?) present an overview of both the possibilities (detection methods) and impossibilities (evasion methods) of AIGT detection. ? (?) focus more specifically on the tasks of author attribution and author obfuscation involving multiple human or AI authors and their combination. ? (?) offer an accessible summary of major insights and challenges in the field of AIGT detection. Two recent pre-print surveys and their associated github pages are also useful resources: ? (?) [^1] https://github.com/Xianjun-Yang/Awesome_papers_on_LLMs_detection and ? (?) [^2] https://github.com/NLP2CT/LLM-generated-Text-Detection. In the current work, we hope to build on existing research to achieve the following goals: (1) To summarize emerging and state-of-the-art methods in a rapidly changing field; and (2) To provide a more practical perspective on the problem, by examining the factors that contribute to detectability in any specific use case, taking into account that not all AIGT is equivalent, and that users of detection software are often working with incomplete knowledge of the model that generated the text, the prompt that was input to the model, the underlying decoding strategy, and any steps that may have been taken to obfuscate the provenance of the text. As such, we hope to produce an approachable and useful document for AI practitioners as well as researchers, by highlighting how different features of AI-generated text can impact the performance of different detection algorithms.
@@ 2.1 Text Classification and Generative AI
In most cases, AIGT detection is approached using a supervised learning framework, which assumes that labelled examples are available to train or calibrate the classifier. As we will see in the following sections, some detectors are trained using the classical approach of first extracting relevant features from the text (e.g. syntactic or stylistic features) and then feeding those features to a machine learning classifier, either statistical (logistic regression, SVM) or neural (deep neural network). Other approaches leverage pre-trained language models (e.g. BERT, ? (?)), which are pre-trained in an unsupervised fashion on large text corpora to learn effective representations of the semantic meaning of text as a series of dense distributions. These models can then be fine-tuned for any number of NLP tasks (including AIGT detection), without the need for an explicit feature extraction step.
While pretrained lagnuage models like BERT are trained using a bidirectional framework to predict a masked token, a generative language model is trained to predict the next word in a sequence, conditioned on the previous words (i.e., the context). The early iterations of generative language models (e.g., GPT by ? (?)) can be thought of simply as pre-trained language models that can generate a full continuation of a text by iteratively generating the next word, as shown in Fig. 2. Such models have become known as large language models, or LLMs.
@@
Nonetheless, the basic functionality of the LLM is still to produce a sequence of words, given the context. At each step in the generation process, the LLM has a probability distribution over all words in its vocabulary. The specific process by which a word is selected from that probability distribution is called the decoding method. A greedy decoding strategy always selects the highest probability word, but this produces deterministic and repetitive generations. In contrast, if the model simply samples the full vocabulary according to the distribution, a poor word choice is likely due to the combined probability mass of all low-probability words. Therefore, a method is needed to choose words that have a high probability given the context (to ensure fluency and consistency) while also sampling from a relatively wide range (to ensure variety and creativity). In practice, two of the most popular decoding strategies are nucleus sampling and top-k sampling; words are sampled according to their probability values, but the choices are still restricted to only the most probable words. Top-$k$ sampling always selects among the top $k$ most probable choices, whereas nucleus sampling uses the total probability sum as the cut-off criterion, meaning that more choices are possible when the model is uncertain. In Section 5.1, we discuss how the decoding strategy can affect the detectability of the generated text.
Both the probability values of the selected words and the shape of the probability distribution at each generation step offer important information about how “surprising” the text is, or how uncertain the model is (refer to Fig. 2). In the NLP literature, there are two commonly used metrics to quantify this uncertainty. Perplexity is related to the inverse of the length-normalized word sequence probability. For example, if every word in a sentence was predicted with a high probability, the sentence is not very surprising, and it has low perplexity (e.g., Fig. 2(a)). Entropy is a measure of uncertainty or spread in the probability distribution. Highly uniform distributions have high entropy and imply that the model is uncertain about the next word given the context (e.g., Fig. 2(b)). In Section 3.2.1, we will see how these statistical properties of language generation can be used to detect AIGT. Because LLMs generate text by always selecting the next word from the set of highest-probability words, AIGT tends to have low perplexity and low entropy, compared to human texts, as measured by the generating model’s probability distribution.
@@ 2.2 Taxonomy of AI-Generated Text
Even more human control is exercised in controlled generation. Here, the complete content of the text is specified, but the language model is used to modify the text in some way: paraphrase, change the style (perhaps make it more professional, or more casual), summarize/shorten, or translate to a different language. This method of producing texts can be extremely powerful, but necessitates the existence of an input text (presumably written by a human, but possibly also itself AI generated). On the boundary between the controlled category and the next category we have “polishing,” a use case in which a human author inputs their own writing but uses AI to make minor changes, e.g. to improve fluency or readability. This use of AI is still under debate in academic and professional settings, as AI proof-reading tools can help writers improve the communication of their ideas in their second language, but may also be detected by plagiarism/AIGT-detection systems (?).
@@ 3.1 Watermarking
A watermark is an identifier that is secretly embedded in text and used to convey some meta-information. In the context of AIGT detection, the idea is that the presence of a watermark denotes that the text is AI-generated. With knowledge of the watermark extraction algorithm, one can verify the text authorship. In other contexts, multi-bit watermarking (?) may be used to encode additional copyright information. Ideally, inserting a watermark does not alter the meaning or quality of the original text, and it should be imperceptible to those without knowledge of the watermarking method. If an adversary detects the use of a watermark, they might change text generation methods or attempt to erase it. A good watermark should therefore be robust to adversarial attacks such as text perturbations and paraphrasing.
? (?) introduce the idea of embedding a watermark during text generation by upsampling specific words in the vocabulary. Specifically, at every token location, the vocabulary is divided into a green list and a red list using a random seed generated from a hash on the preceding token. The probability distribution is biased to select green words so that they appear with greater frequency in watermarked text. With knowledge of the hash function, we can observe the ratio of green to red words and detect the watermark using a statistical hypothesis test. As the text sequence length increases, the probability of misclassifying a text as a false positive is vanishingly small. However, the structured nature of using the preceding token in computing the red-green list introduces vulnerability to perturbation attacks (i.e., making minor modifications or perturbations to the text). Word deletions, insertions, and synonym replacements can all weaken the watermark signal. As an extreme example, substituting every other word in the passage with a synonym will destroy the watermark completely. Later ? (?) simplified the idea to use a fixed red-green list which was shown to improve robustness to perturbation attacks. The trade-off for enhanced robustness is that an adversary can more easily detect and learn the red-green list, and hence remove or forge the watermark. Note however that an adversary could reconstruct the scheme in either case, given unlimited (black-box) access to the watermark-generating LLM.
@@
For any closed-source model, white-box methods as described above leave all responsibility for watermarking to the owners. In contrast, black-box methods take only the generated text as input, enabling anyone to inject a watermark post-generation. This can be useful for third parties that build applications on top of closed-source language models and want to indicate that text is AI-generated or otherwise embed some copyright information. Early work included formatting strategies such as line shifts (?) and Unicode character replacement (?), however, these are easily detected and can be removed through text canonicalization (?). Other methods encode information through syntactic structures (?) or vocabulary choice using synonym replacement similar to the idea of red-green lists (?). In the black box setting, synonym replacement is prone to degrade text quality because the substitution is not selected using the LM’s probability distribution. ? (?) improves the text quality of black-box lexical-based watermarking by using BERT to select context-aware synonyms. Note that many simple watermark techniques are complementary and therefore can be stacked together. For instance, both lexical and syntactic watermarks can be applied simultaneously. The syntactic watermark provides some robustness to synonym replacement attacks that lexical watermarks are vulnerable to. However, both strategies are vulnerable to paraphrase attacks that alter both syntax and word choice. One counter idea is to force the adversary to change the most important features of the text (e.g. proper nouns, semantically essential words) to remove the watermark (?), although the quality of watermarked text is also compromised by this approach. In general, robust black-box watermarking methods seem to be lacking. A workaround could be to implement white-box methods post-generation by taking the text to watermark, asking an open-source model for a paraphrase, and applying the watermark during the paraphrase generation. Of course, this strategy is only attractive if an open-source model can provide a regeneration of similar quality to the original closed-source generation. ? (?) introduces an end-to-end regeneration strategy (Remark-LLM) that first integrates the watermark into the learned semantic representation, and then uses a trained decoder module to extract the watermark. By using a beam search in the decoding step, the regeneration is optimized for coherence and consistency. Because Remark-LLM is trained end-to-end, the authors include malicious examples in the training data to improve robustness. Compared to rule-based approaches, neural watermark embedding requires more text to encode the same amount of information.
@@ 3.2.1 Statistical Analysis
The most trivial implementation of a statistical classifier is to directly measure the likelihood of generating some observed text according to a particular model of interest. For each word in a passage, we can use the language model to rank the word choice as well as compute its probability value. Averaging the token-wise probability or rank over the text passage gives us an interpretable score for the text, where higher probabilities and lower ranks indicate AI generation. ? (?) propose using the ratio of average probability and rank since they provide complementary information about the model’s certainty of the word choice. The entropy of a text span can also be computed from the probability values to give us an idea of the model’s certainty in a word choice given the context. In all cases, these measures are designed to tell us how likely it is that the text was generated by a known model, and some threshold value can be set for classification. One advantage of word-level measures is that they allow for intuitive visualization tools (GLTR, ? (?)) that can help humans understand and recognize AIGT. Another advantage is that they allow for flexible rolling averages over the text passage (i.e. we can detect how measures evolve over a document or concatenation of time-ordered social media posts).
@@
While word-level measures look at the value of the language model’s probability function in various ways, we can also learn something from the structure of the probability function. ? (?) hypothesize and empirically demonstrate that AIGT tends to occupy the negative curvature regions of the probability function, meaning that any small perturbations of the text will result in lower probability values. Leveraging this observation, they develop DetectGPT, a single-feature method based on a simple computation of the average change in log probability over a sample of text perturbations. While DetectGPT was designed for the white-box, known-model setting, an assumption that we have access to the model that we are trying to detect, experimental results (?, ?) indicate some transferability to the unknown-model setting. ? (?) demonstrate that cross-detection can perform almost as well as self-detection, and notably, smaller language models are better universal detectors. Also based on the susceptibility of AIGT to local perturbations, ? (?) define a single-feature measure for the average perturbed rank (NPR). Experimentally, both DetectGPT and NPR achieve impressive detection of GPT2-era generation models in the domain of news and prompted stories. More recently, ? (?) propose an update to DetectGPT called Fast-DetectGPT, which improves the efficiency of DetectGPT while also increasing the accuracy: experimental results on text from ChatGPT and GPT-4 show Fast-DetectGPT outperforming both DetectGPT and NPR. Finally, ? (?) observe that language models tend to transmit information more uniformly than humans, and therefore suggest the variance in word probability over a text span as an informative feature for classification.
@@
For many practical scenarios, these methods are limited because we cannot access the probability values of closed-source models (e.g. GPT-4 at the time of writing). In this case, one way to determine whether text was generated from a particular model is to simply regenerate the text and compare the resulting generation to the original. In this setup, the first portion of the text is retained as the input to the language model, and we do not assume knowledge of the actual prompt that was used. ? (?) provide both black-box and white-box variations of this idea, where they compare either output words or probability values, respectively. They demonstrate superior detection performance compared to DetectGPT in the white-box setting, and language model-based strategies in the black-box setting. Furthermore, they demonstrate robustness to AI-editing (i.e. perturbations made by another language model). However, the results do depend heavily on the number of regenerations and the truncation ratio (i.e., how much text is used for the input prompt versus the output comparison), and therefore some domain-specific development data may be needed for hyper-parameter selection. More specific to the instruction interface of ChatGPT, ? (?) propose a regeneration method in which ChatGPT both generates the prompt from the original text and then regenerates the output. In contrast to language model-based detection methods, the regeneration strategies claim better generalization to unseen data domains and emerging language models, and better robustness to word substitutions, re-translation, and hybrid human-AI texts. Note that these methods are most suited to arbitrary or guided generation (i.e. prompts that are not overly specific) in the black-box setting, and it is unclear if they would be successful in detecting controlled generation.

### v2

@@ 1 Introduction
In recent years, the capabilities of large language models (LLMs) to generate fluent, realistic-sounding text have improved dramatically. We are now at a point where humans themselves cannot reliably distinguish between text that was generated using artificial intelligence (AI) and that written by a real person (?, ?, ?). There are many opportunities for LLMs to contribute to human productivity, with applications to question-answering, computer programming, brainstorming, proof-reading, and information retrieval. However, LLMs can also facilitate malicious activities, increasing efficiency and reducing costs in malware creation, fraud, identity theft, harassment attacks, and academic dishonesty (?, ?) (see Figure 1). Another major risk brought by generative AI is the potential for global-scale “information pollution.” Automatically generated ‘fake’ texts can be exploited for commercial goals, such as product promotion or fake product reviews, or for political gains, encompassing propaganda, ‘fake news’, disinformation, etc. Current state-of-the-art generative models can produce high-quality, fluent fake information that is perceived as more credible and trustworthy than human-generated misinformation (?, ?), and that is harder for both human readers and automatic detection systems to recognize (?, ?, ?).
Figure 1: Some common types of misuse of AI-generated text (?, ?).
@@
The present manuscript provides an overview of the field of AIGT detection, the current state-of-the-art methodologies, the available data resources and online tools, and existing challenges. In this survey, we have included papers published in both peer-reviewed journals and conference proceedings, as well as those posted on pre-print servers such as arXiv. While pre-print manuscripts may not have been formally peer-reviewed yet, we include them in an attempt to cover the most cutting-edge advancements in a rapidly changing field. We have focused our attention primarily on papers published in 2023 and 2024, with some references to seminal papers published earlier. [^1] The cut-off date for this survey was 10 June, 2024. However, we have updated the reference information for papers that appeared as pre-prints before the cut-off and were subsequently published. We include only those papers that focus on NLP methods for AIGT detection, excluding methods that rely on other characteristics of bot accounts, such as posting frequency or social network analysis. However, within the realm of NLP, we cast our net widely, including methods from diverse data domains (social media, news stories, academic essays, scientific abstracts) and aiming to provide an extensive and high-level overview of the field as it currently stands.
@@
Other surveys on AIGT detection exist, and we point the interested reader to them for more information. ? (?) and ? (?) focus specifically on watermarking for LLMs. ? (?) provide an extensive analysis of the risks of AI-generated text, and an evaluation of methods to detect text generated by pre-ChatGPT models. ? (?) present an overview of both the possibilities (detection methods) and impossibilities (evasion methods) of AIGT detection. ? (?) focus more specifically on the tasks of author attribution and author obfuscation involving multiple human or AI authors and their combination. ? (?) offer an accessible summary of major insights and challenges in the field of AIGT detection. Two recent surveys and their associated github pages are also useful resources: ? (?) [^2] https://github.com/Xianjun-Yang/Awesome_papers_on_LLMs_detection and ? (?) [^3] https://github.com/NLP2CT/LLM-generated-Text-Detection. In the current work, we hope to build on existing research to achieve the following goals: (1) To summarize emerging and state-of-the-art methods in a rapidly changing field; and (2) To provide a more practical perspective on the problem, by examining the factors that contribute to detectability in any specific use case, taking into account that not all AIGT is equivalent, and that users of detection software are often working with incomplete knowledge of the model that generated the text, the prompt that was input to the model, the underlying decoding strategy, and any steps that may have been taken to obfuscate the provenance of the text. As such, we hope to produce an approachable and useful document for AI practitioners as well as researchers, by highlighting how different features of AI-generated text can impact the performance of different detection algorithms.
@@ 2.1 Text Classification and Generative AI
In most cases, AIGT detection is approached using a supervised learning framework, which assumes that labelled examples are available to train or calibrate the classifier. As we will see in the following sections, some detectors are trained using the classical approach of first extracting relevant features from the text (e.g., syntactic or stylistic features) and then feeding those features to a machine learning classifier, either statistical (logistic regression, SVM) or neural (deep neural network). Other approaches leverage pre-trained language models (e.g., BERT, ? (?)), which are pre-trained in an unsupervised fashion on large text corpora to learn effective representations of the semantic meaning of text as a series of dense distributions. These models can then be fine-tuned for any number of NLP tasks (including AIGT detection), without the need for an explicit feature extraction step.
While pretrained language models like BERT are trained using a bidirectional framework to predict a masked token, a generative language model is trained to predict the next word in a sequence, conditioned on the previous words (i.e., the context). The early iterations of generative language models (e.g., GPT by ? (?)) can be thought of simply as pre-trained language models that can generate a full continuation of a text by iteratively generating the next word, as shown in Figure 2. Such models have become known as large language models, or LLMs.
@@
Nonetheless, the basic functionality of the LLM is still to produce a sequence of words, given the context. At each step in the generation process, the LLM has a probability distribution over all words in its vocabulary. The specific process by which a word is selected from that probability distribution is called the decoding method. A greedy decoding strategy always selects the highest probability word, but this produces deterministic and repetitive generations. In contrast, if the model simply samples the full vocabulary according to the distribution, a poor word choice is likely due to the combined probability mass of all low-probability words. Therefore, a method is needed to choose words that have a high probability given the context (to ensure fluency and consistency) while also sampling from a relatively wide range (to ensure variety and creativity). In practice, two of the most popular decoding strategies are nucleus sampling and top-k sampling; words are sampled according to their probability values, but the choices are still restricted to only the most probable words. Top-$k$ sampling always selects among the top $k$ most probable choices, whereas nucleus sampling uses the total probability sum as the cut-off criterion, meaning that more choices are possible when the model is uncertain (see Figure 2). In Section 5.1, we discuss how the decoding strategy can affect the detectability of the generated text.
Both the probability values of the selected words and the shape of the probability distribution at each generation step offer important information about how “surprising” the text is, or how uncertain the model is (refer to Figure 2). In the NLP literature, there are two commonly used metrics to quantify this uncertainty. Perplexity is related to the inverse of the length-normalized word sequence probability. For example, if every word in a sentence was predicted with a high probability, the sentence is not very surprising, and it has low perplexity (e.g., Figure 2(a)). Entropy is a measure of uncertainty or spread in the probability distribution. Highly uniform distributions have high entropy and imply that the model is uncertain about the next word given the context (e.g., Figure 2(b)). In Section 3.2.1, we will see how these statistical properties of language generation can be used to detect AIGT. Because LLMs generate text by always selecting the next word from the set of highest-probability words, AIGT tends to have low perplexity and low entropy, compared to human texts, as measured by the generating model’s probability distribution.
@@ 2.2 Taxonomy of AI-Generated Text
Even more human control is exercised in controlled generation. Here, the complete content of the text is specified, but the language model is used to modify the text in some way: paraphrase, change the style (perhaps make it more professional, or more casual), summarize/shorten, or translate to a different language. This method of producing texts can be extremely powerful, but necessitates the existence of an input text (presumably written by a human, but possibly also itself AI generated). On the boundary between the controlled category and the next category we have “polishing,” a use case in which a human author inputs their own writing but uses AI to make minor changes, e.g., to improve fluency or readability. This use of AI is still under debate in academic and professional settings, as AI proof-reading tools can help writers improve the communication of their ideas in their second language, but may also be detected by plagiarism/AIGT-detection systems (?).
@@ 3.1 Watermarking
Figure 3: Generating text with watermarking using red-green lists (?).
A watermark is an identifier that is secretly embedded in text and used to convey some meta-information. In the context of AIGT detection, the idea is that the presence of a watermark denotes that the text is AI-generated. With knowledge of the watermark extraction algorithm, one can verify the text authorship. In other contexts, multi-bit watermarking (?, ?) may be used to encode additional copyright information. Ideally, inserting a watermark does not alter the meaning or quality of the original text, and it should be imperceptible to those without knowledge of the watermarking method. If an adversary detects the use of a watermark, they might change text generation methods or attempt to erase it. A good watermark should therefore be robust to adversarial attacks such as text perturbations and paraphrasing.
? (?) introduce the idea of embedding a watermark during text generation by upsampling specific words in the vocabulary (Figure 3). Specifically, at every token location, the vocabulary is divided into a green list and a red list using a random seed generated from a hash on the preceding token. The probability distribution is biased to select green words so that they appear with greater frequency in watermarked text. With knowledge of the hash function, we can observe the ratio of green to red words and detect the watermark using a statistical hypothesis test. As the text sequence length increases, the probability of misclassifying a text as a false positive is vanishingly small. However, the structured nature of using the preceding token in computing the red-green list introduces vulnerability to perturbation attacks (i.e., making minor modifications or perturbations to the text). Word deletions, insertions, and synonym replacements can all weaken the watermark signal. As an extreme example, substituting every other word in the passage with a synonym will destroy the watermark completely. Later ? (?) simplified the idea to use a fixed red-green list which was shown to improve robustness to perturbation attacks. The trade-off for enhanced robustness is that an adversary can more easily detect and learn the red-green list, and hence remove or forge the watermark. Note however that an adversary could reconstruct the scheme in either case, given unlimited (black-box) access to the watermark-generating LLM.
@@
For any closed-source model, white-box methods as described above leave all responsibility for watermarking to the owners. In contrast, black-box methods take only the generated text as input, enabling anyone to inject a watermark post-generation. This can be useful for third parties that build applications on top of closed-source language models and want to indicate that text is AI-generated or otherwise embed some copyright information. Early work included formatting strategies such as line shifts (?) and Unicode character replacement (?), however, these are easily detected and can be removed through text canonicalization (?). Other methods encode information through syntactic structures (?) or vocabulary choice using synonym replacement similar to the idea of red-green lists (?). In the black box setting, synonym replacement is prone to degrade text quality because the substitution is not selected using the LM’s probability distribution. ? (?) improves the text quality of black-box lexical-based watermarking by using BERT to select context-aware synonyms. Note that many simple watermark techniques are complementary and therefore can be stacked together. For instance, both lexical and syntactic watermarks can be applied simultaneously. The syntactic watermark provides some robustness to synonym replacement attacks that lexical watermarks are vulnerable to. However, both strategies are vulnerable to paraphrase attacks that alter both syntax and word choice. One counter idea is to force the adversary to change the most important features of the text (e.g., proper nouns, semantically essential words) to remove the watermark (?), although the quality of watermarked text is also compromised by this approach. In general, robust black-box watermarking methods seem to be lacking. A workaround could be to implement white-box methods post-generation by taking the text to watermark, asking an open-source model for a paraphrase, and applying the watermark during the paraphrase generation. Of course, this strategy is only attractive if an open-source model can provide a regeneration of similar quality to the original closed-source generation. ? (?) introduces an end-to-end regeneration strategy (Remark-LLM) that first integrates the watermark into the learned semantic representation, and then uses a trained decoder module to extract the watermark. By using a beam search in the decoding step, the regeneration is optimized for coherence and consistency. Because Remark-LLM is trained end-to-end, the authors include malicious examples in the training data to improve robustness. Compared to rule-based approaches, neural watermark embedding requires more text to encode the same amount of information.
@@ 3.2 Statistical and Stylistic Analysis
Both statistical and stylistic analysis involve the extraction of features from the generated text. Statistical methods measure how probable, or likely, each word in the sequence is (relative to some probability distribution). Stylistic methods focus on linguistic properties of the generated text, such as vocabulary, syntax, and coherence.
@@ 3.2.1 Statistical Analysis
The most trivial implementation of a statistical classifier is to directly measure the likelihood of generating some observed text according to a particular model of interest. For each word in a passage, we can use the language model to rank the word choice as well as compute its probability value. Averaging the token-wise probability or rank over the text passage gives us an interpretable score for the text, where higher probabilities and lower ranks indicate AI generation. ? (?) propose using the ratio of average probability and rank since they provide complementary information about the model’s certainty of the word choice. The entropy of a text span can also be computed from the probability values to give us an idea of the model’s certainty in a word choice given the context. In all cases, these measures are designed to tell us how likely it is that the text was generated by a known model, and some threshold value can be set for classification. One advantage of word-level measures is that they allow for intuitive visualization tools (GLTR, ? (?)) that can help humans understand and recognize AIGT. Another advantage is that they allow for flexible rolling averages over the text passage (i.e., we can detect how measures evolve over a document or concatenation of time-ordered social media posts).
Figure 4: AIGT tends to occupy the negative curvature regions of the probability function (?).
@@
While word-level measures look at the value of the language model’s probability function in various ways, we can also learn something from the structure of the probability function. ? (?) hypothesize and empirically demonstrate that AIGT tends to occupy the negative curvature regions of the probability function, meaning that any small perturbations of the text will result in lower probability values (see Figure 4). Leveraging this observation, they develop DetectGPT, a single-feature method based on a simple computation of the average change in log probability over a sample of text perturbations. While DetectGPT was designed for the white-box, known-model setting, an assumption that we have access to the model that we are trying to detect, experimental results (?, ?) indicate some transferability to the unknown-model setting. ? (?) demonstrate that cross-detection can perform almost as well as self-detection, and notably, smaller language models are better universal detectors. Also based on the susceptibility of AIGT to local perturbations, ? (?) define a single-feature measure for the average perturbed rank (NPR). Experimentally, both DetectGPT and NPR achieve impressive detection of GPT2-era generation models in the domain of news and prompted stories. More recently, ? (?) propose an update to DetectGPT called Fast-DetectGPT, which improves the efficiency of DetectGPT while also increasing the accuracy: experimental results on text from ChatGPT and GPT-4 show Fast-DetectGPT outperforming both DetectGPT and NPR. Finally, ? (?) observe that language models tend to transmit information more uniformly than humans, and therefore suggest the variance in word probability over a text span as an informative feature for classification.
@@
For many practical scenarios, these methods are limited because we cannot access the probability values of closed-source models (e.g., GPT-4 at the time of writing). In this case, one way to determine whether text was generated from a particular model is to simply regenerate the text and compare the resulting generation to the original. In this setup, the first portion of the text is retained as the input to the language model, and we do not assume knowledge of the actual prompt that was used. ? (?) provide both black-box and white-box variations of this idea, where they compare either output words or probability values, respectively. They demonstrate superior detection performance compared to DetectGPT in the white-box setting, and language model-based strategies in the black-box setting. Furthermore, they demonstrate robustness to AI-editing (i.e., perturbations made by another language model). However, the results do depend heavily on the number of regenerations and the truncation ratio (i.e., how much text is used for the input prompt versus the output comparison), and therefore some domain-specific development data may be needed for hyper-parameter selection. More specific to the instruction interface of ChatGPT, ? (?) propose a regeneration method in which ChatGPT both generates the prompt from the original text and then regenerates the output. In contrast to language model-based detection methods, the regeneration strategies claim better generalization to unseen data domains and emerging language models, and better robustness to word substitutions, re-translation, and hybrid human-AI texts. Note that these methods are most suited to arbitrary or guided generation (i.e., prompts that are not overly specific) in the black-box setting, and it is unclear if they would be successful in detecting controlled generation.